how to do square root of 2 plus square root of 1/2
Found 2 solutions by solver91311, RAY100:
Answer by solver91311(24713)





First, apply





Next, rationalize your denominator by multiplying by 1 in the form of



Now apply the LCD of 2:



John


Answer by RAY100(1637)

sq rt 2 + sq rt 1/2 = 2^ (1/2) + (1/2) ^ (1/2) = 1.141 + ,707 = 2.121